# Extending the SAREF4ENER Ontology with Flexibility Based on FlexOffers

### v1

@@ Abstract
A key element to support the increased amounts of renewable energy in the energy system is flexibility, i.e., the possibility of changing energy loads in time and amount. Many flexibility models have been designed; however, exact models fail to scale for long time horizons or many devices. Because of this, the FlexOffer (FOs) model has been designed, to provide device-independent approximations of flexibility with good accuracy, and much better scaling for long time horizons and many devices. An important aspect of the real-life implementation of energy flexibility is enabling flexible data exchange with many types of smart energy appliances and market systems, e.g., in smart buildings. For this, ontologies standardizing data formats are required. However, the current industry standard ontology for integrating smart devices for energy purposes, SAREF for Energy Flexibility (SAREF4ENER) only has limited support for flexibility and thus cannot support important use cases. In this paper we propose an extension of SAREF4ENER that integrates full support for the complete FlexOffer model, including advanced use cases, while maintaining backward compatibility. This novel ontology module can accurately describe flexibility for advanced devices such as electric vehicles, batteries, and heat pumps. It can also capture the inherent uncertainty associated with many flexible load types.
@@ I Introduction
The last decades have seen an increasing usage of renewable energy, and the development of new technologies and concepts to enable its exploitation. In particular, among those new concepts, one has crucial importance for handling the intermittency that comes with renewable energy generation: flexibility [1], i.e., the capability of changing energy loads in time or amount of energy. Flexibility enables optimizing energy consumption to achieve several objectives: renewable energy usage maximization [2], peak shaving [3, 4], CO2 emissions minimization [5], energy cost minimization [6, 7], provision of ancillary services [8], congestion prevention [9], and demand response [10]. Several models have been developed for flexibility; many of these models are accurate in their description of flexibility but scale poorly with respect to the length of time horizons and the number of devices. A model that solves these issues is FlexOffers (FOs) [11], a device-independent model that represents flexibility by using energy constraints that approximate the amount of available flexibility, but scales very well in scenarios with many devices or long time horizons. FOs is a powerful model used in more than 30 EU and national projects, whose real-life implementation has been advancing in the last decade. An important aspect is its interaction with smart devices/buildings and the ontologies representing them. To date, a widely used ontology for smart buildings is the Smart Application REFerence (SAREF) [12] ontology. Specifically, several extensions of SAREF have been created for different domains: SAREF for Energy (SAREF4ENER) [13] has been designed to represent energy devices and services, and is the de-facto standard for that. However, SAREF4ENER fails to capture energy flexibility for many use cases, specifically when flexibility constraints become more complex: this is the case for advanced devices such as batteries, electric vehicles (EVs), and heat pumps. This raises the need for an ontology capable of handling these cases. To meet this challenge, an ontology called domOS Common Ontology (dCO  [^1] https://w3id.org/dco[14]has been proposed, to ensure full flexibility support in the interoperability of smart buildings and services. The contribution we bring in this paper is the dCO flexibility module: this module extends SAREF4ENER in its representation of flexibility services, enabling compatibility with the full FO model and constraints and ensuring that the most relevant flexible devices and use cases for them are represented. In this paper, we will show how the flexibility module of dCO extends SAREF4ENER, how it represents the full FO model and constraints, and some advanced use cases that can be represented by it. The remainder of this paper is organized as follows. Section II describes FOs and the SAREF4ENER and dCO ontologies, while Section III shows how dCO extends SAREF4ENER and enables full FO representation and Section IVevaluates the dCO flexibility module, showing which use cases can and cannot be represented by FOs and SAREF4ENER. Finally, Section V concludes the work and provides ideas for future research.
@@ II-A Running Example
We start by making a running example so that we can describe the FO concept in the following subsections. Here, we will consider the case of a heat pump operating in a room. The heat pump is a Daikin Altherma [15], with maximum operational power $P_{max}=4.6$ kW and coefficient of performance (COP) $3.65$. This model supports the SG-Ready interface [16, 17]. The room has a rectangular section with sides of $4$ and $5$ meters respectively, and it is $3$ meters high. The heat transfer coefficient is $c_{ht}=6\frac{W}{m^{2}K}$ [18], and heat is dispersed through one of the $3$m x $4$m side walls. The room temperature is $295$K ($22^{\circ}$), while the outside temperature is $275$K ($2^{\circ}$). We assume that room temperature must stay between $293$K ($20^{\circ}$) and $297$K ($24^{\circ}$) respectively, because of user settings. For reference, the amount of energy (heat) needed to maintain the temperature to $295$K for one hour can be obtained by
@@ II-B FlexOffer Life Cycle
The FO model allows for describing flexibility in a unified format for different device types [11, 19], by providing an approximated representation of flexibility. In exchange for the small amount of accuracy lost with respect to exact models, modeling flexibility with FOs allows for much better scalability over long time horizons and high numbers of devices. FOs provide constraints for energy consumption: optimizing an FO means to solve an optimization problem on energy consumption (e.g., cost minimization) within the constraints given by the FO. The solution of this optimization problem is a set of values $(e_{1},\ldots,e_{T})$ for energy consumption at each time unit: this is called a schedule. FOs can also be aggregated: this means that, from $M$ FOs, it is possible to generate $N\ll M$ FOs that represent their combined flexibility, with some losses. The opposite process is called disaggregation: from a schedule for an aggregated FO, it is possible to generate schedules for the original FOs so that their combined energy consumption is the same as the original schedule.
@@
Figure 1: A schematic description of the FO life cycle (taken from [21]).
@@ II-C Standard FlexOffers
The simplest FO constraints are start time constraints and slice (energy) constraints. We are considering time as discrete and divided in time units, also called time slices: their duration is one hour in the running example. Start time constraints define the earliest and latest time at which the load can start. Energy constraints define, at each time unit, the minimum and maximum amount of energy that can be consumed from the considered energy load: in other words, at each time unit $t$, two numbers $emin_{t}$ and $emax_{t}$ are defined such that $emin_{t}\leq e_{t}\leq emax_{t}$, where $e_{t}$ is the energy consumption during time unit $t$. An FO with just these constraints is called a Standard FlexOffer (SFO). We can see an SFO example in Figure 2(a) for the running example for eight time units. The pink bars indicate the amount of energy that can be consumed at each time unit: here, at each time unit, it varies between 0.303 kWh and 0.478 kWh. At each time, the horizontal bars indicate a possible amount of energy consumption $e_{t}$.
@@ II-F Uncertain FlexOffers
Uncertain FOs (UFOs) [21] are a type of FOs that consider uncertainty from flexibility devices behavior. Two main types of uncertainty are considered: time and amount of uncertainty. At each time unit, time uncertainty refers to the probability for the device to be operative at that time, while amount uncertainty refers to the amount of energy that can be consumed by the device at that time. A UFO is created in two steps. First, uncertainty related to the state (in the running example, temperature) is modeled at each time $t$; second, the probability for each energy value at each time to be feasible is calculated, taking into account both types of uncertainty. Some functions $\{f_{1},\ldots,f_{T}\}$ describing those probabilities will be obtained: those functions will define the UFO. UFOs can be visualized by choosing a probability threshold $P_{0}$. At each time $t$, the energy values having probability at least $P_{0}$ of being available for consumption can be described by intervals. Figure 2(d) shows this visualization for the running example: if we choose $P_{0}=1$, at each time unit, the available energy values are described by the pink bars, while if we choose $P_{0}=0.8$, the available energy values are described by the combined pink and blue bars. This is because at the first time unit, the status of the device is known, and the amount of available flexibility can be determined surely, while for the next time units, there is more uncertainty (represented here by the blue bars) and the amount of flexibility certainly available (pink bars) is lower.
@@ II-G SAREF and SAREF4ENER
Smart Applications REFerence (SAREF) [^2] https://saref.etsi.org/ [12] is an ontology for IoT devices and solutions published by ETSI [^3] https://www.etsi.org/ in a series of Technical Specifications initially released in 2015 and updated in 2017. Even if its initial objective was to build a reference ontology for appliances relevant to energy efficiency, SAREF can also serve as a reference model to enable better integration of data from various vertical domains in the IoT. SAREF has been extended to different domains: specifically, SAREF for Energy Flexibility (SAREF4ENER) [13] is its extension for the energy domain, and aims to provide a standardized way to describe and interoperate energy-related devices, systems, and services. It allows us to describe devices such as solar panels, electric vehicles, and smart meters, as well as systems like microgrids, and energy management/demand response services. By using SAREF4ENER, developers can create applications that easily integrate and communicate with different energy devices, systems, and services, regardless of the manufacturer or protocol used, which enables more efficient and sustainable energy management.
@@
The World Wide Web Consortium (W3C) has developed a standard for enabling Internet of Things (IoT) interoperability, whose name is W3C Web of Things (WoT) [22]. In this context, Thing Descriptions (TDs) [23] have been developed: they describe the metadata and interfaces of the considered devices that participate in the WoT. Things are organized in TD Directories (WoT TDDs), which make it possible to register, manage, and search TD databases. An important feature for IoT consumers is to send queries to the WoT directory: this is achieved by having a common ontology to annotate TDs; however, before dCO, no ontologies were integrating WoT into smart buildings. dCO [14] has been created with the purpose of achieving this integration: in this paper, we will focus on the aspect of dCO that strictly concerns energy flexibility. A description of the modules can be found in Appendix -A). The contribution of our work is the energy flexibility module, which is not described in [14]: in this paper, we explain how it extends SAREF4ENER.
@@
This section shows how SAREF4ENER has been extended to describe FOs. We show how we designed the dCO flexibility module and how we encode attributes and constraints. Details on FO message can be found in Appendix -B.
@@ III-B Energy Flexibility Ontological Design
  • hasEnergyConstraintList: (max 1 s4ener:EnergyMin, max 1 s4ener:EnergyMax. In Figure 2(a), they are 0.303 and 0.478 respectively);
@@
  • dco:totalEnergyConstraint: (max 1 s4ener:EnergyMin, max 1 s4ener:EnergyMax. In Figure 2(b), they are 2.592 and 3.381 respectively);
  • dependencyEnergyConstraintList: Contains the dependency constraints for energy in a Dependency FlexOffer;
  • uncertainFunctions: (min 1 Function): Contains the parameters which define the uncertain FlexOffer;
  • uncertainThreshold: a real number that represents the probability threshold. In the running example in Figure 2(d), it can be 1 (pink) or 0.8 (blue).
@@
  • scheduleID: the id of a schedule;
  • updateID: the id of the schedule update;
@@ IV Evaluation
The project demonstrated dCO feasibility in real-world case studies in Denmark, France and Switzerland: all functionalities, e.g., querying building descriptions, are operative, and interoperability between smart buildings and services achieved. The energy services using dCO cut down energy consumption and smoothened load consumption curves for 124 individual customers and 50 days. The cumulative amount of flexibility (load shedding) was several mWh.
Regarding flexibility representation, SAREF4ENER is able at most to accurately represent flexibility for devices that only need this type of constraint. This is the case with wet devices, such as dishwashers and washing machines, as shown in [11]. These devices have a load with a fixed power/amount profile, which is only shifted in time. However, for other devices, SFOs/TECFOs are outclassed by DFOs in terms of accuracy in flexibility representation, as we can see in previous literature. The ability of retaining flexibility for a model is calculated by the economic metric [19], defined as the amount of profit that can be obtained by optimizing flexibility with the objective of cost reduction. Specifically, for the charge and discharge process of batteries, DFOs retain $61$% of the flexibility, against the $38$% from TECFOs and $10$% from SFOs [19]. UFOs can perform even better, retaining up to $66.4$% of flexibility, which goes up to $86.8$% of flexibility when imbalance penalties for violating constraints are high (high penalty case) [21]. Regarding EVs, flexibility modeling via DFOs is more accurate than SFOs and TECFOs, as slice constraints and TECs are only a subset of DFO constraints, and therefore the amount of flexibility retained by DFOs is higher: yet, in [21], we see that UFOs can retain $92$% of flexibility, against the $77.3$% retained by DFOs. For the high penalty case, UFOs retain $87.5$% of flexibility, against the $61.8$% from DFOs, which still outperform SFOs and TECFOs. Finally, regarding heat pumps, the comparison between DFOs and SFOs has been done in [24], showing that DFOs can retain more than three times the amount of flexibility from SFOs: in [15] DFOs for heat pumps have been further improved, showing that they can retain up to $98.9$% of flexibility according to the economic metric. The ability for dCO to model the significant amounts of flexibility for these devices ensures that developers can create applications that allow for an accurate description of flexibility for batteries, EVs and heat pumps. Table I resumes this, showing which use cases are supported by the dCO flexibility module and SAREF4ENER, respectively.
@@
In order to validate dCO, we also used the OntOlogy Pitfall Scanner! (OOPS!) tool [^4] https://oops.linkeddata.es/ to evaluate it. OOPS! [25] is a tool, well-known in the Semantic Web community, designed with the purpose of identifying pitfalls and design errors in ontologies. There are three levels of importance in pitfalls identified by OOPS!. Critical pitfalls need to be resolved, as they may affect the ontology consistency, reasoning and applicability. Important pitfalls would improve the ontology’s robustness if fixed, but they do not undermine the ontology function. Minor pitfalls would improve the ontology’s readability if fixed, but do not pose problems for the ontology. OOPS! identified quite a few minor and important pitfalls, but only 11 critical ones. However, the critical pitfalls were all of the type Defining multiple domains or ranges in properties. In our case, it was not only valid, but also necessary for properties to have multiple domains and ranges to capture the complexity of our flexibility scenario. A simple example of this is the dco:hasEnergyConstraintList property, which can have up to one s4ener:EnergyMin attribute and up to one s4ener:EnergyMax attribute. For this reason, we intentionally designed our ontology with this feature, and therefore none of the critical pitfalls identified by the OOPS! tool undermines the function of the ontology, rather they are exactly as they should be. Thus, the ontology meets the checked requirements.

### v2

@@ Abstract
A key element to support the increased amounts of renewable energy in the energy system is flexibility, i.e., the possibility of changing energy loads in time and amount. Many flexibility models have been designed; however, exact models fail to scale for long time horizons or many devices. Because of this, the FlexOffers model has been designed, to provide device-independent approximations of flexibility with good accuracy, and much better scaling for long time horizons and many devices. An important aspect of the real-life implementation of energy flexibility is enabling flexible data exchange with many smart energy appliances and market systems, e.g., in smart buildings. For this, ontologies standardizing data formats are required. However, the current industry standard ontology for integrating smart devices for energy purposes, SAREF for Energy Flexibility (SAREF4ENER), only has limited support for flexibility and thus cannot support important use cases. In this paper, we propose an extension of SAREF4ENER that integrates full support for the complete FlexOffer model, including advanced use cases, while maintaining backward compatibility. This novel ontology module can accurately describe flexibility for advanced devices such as electric vehicles, batteries, and heat pumps. It can also capture the inherent uncertainty associated with many flexible load types.
@@ I Introduction
The last decades have seen an increasing usage of renewable energy, and the development of new technologies and concepts to enable its exploitation. In particular, among those new concepts, one has crucial importance for handling the intermittency that comes with renewable energy generation: flexibility [1], i.e., the capability of changing energy loads in time or amount of energy. Flexibility enables optimizing energy consumption to achieve several objectives: renewable energy usage maximization [2], peak shaving [3, 4], CO2 emissions minimization [5], energy cost minimization [6, 7], provision of ancillary services [8], congestion prevention [9], and demand response [10]. Several models have been developed for flexibility. Many of them are accurate in their description of flexibility, but scale poorly with respect to the length of time horizons and the number of devices. A model that solves these issues is FlexOffers [11], a device-independent model that represents flexibility by using energy constraints that approximate the amount of available flexibility, but scales very well in scenarios with many devices or long time horizons. FlexOffers generates flexoffers (FOs) objects, which describe flexibility from the considered devices.FlexOffers is a powerful model used in more than 30 European and national projects, whose real-life implementation has been advancing in the last decade. An important aspect is its interaction with smart devices/buildings and the ontologies representing them. To date, a widely used ontology for smart buildings is the Smart Application REFerence (SAREF) [12] ontology. Specifically, several extensions of SAREF have been created for different domains: SAREF for Energy (SAREF4ENER) [13] has been designed to represent energy devices and services, and is the de facto standard for that. However, SAREF4ENER fails to capture energy flexibility for many use cases, specifically when flexibility constraints become more complex: this is the case for advanced devices such as batteries, electric vehicles (EVs), and heat pumps. This raises the need for an ontology capable of handling these cases. To meet this challenge, an ontology called domOS Common Ontology (dCO [^1] https://w3id.org/dco[14] has been proposed to ensure full flexibility support in the interoperability of smart buildings and services. The contribution we bring in this paper is the dCO flexibility module: this module extends SAREF4ENER in its representation of flexibility services, enabling compatibility with the full Flexoffers model and constraints, and ensuring that the most relevant flexible devices and use cases for them are represented. In this paper, we show how the flexibility module of dCO extends SAREF4ENER, how it represents the full Flexoffer model and constraints, and some advanced use cases that can be represented by it. The remainder of this paper is organized as follows. Section II describes FOs and the SAREF4ENER and dCO ontologies, while Section III shows how dCO extends SAREF4ENER and enables full FO representation. Section IV evaluates the dCO flexibility module, showing which use cases can and cannot be represented by FOs and SAREF4ENER. Finally, Section V concludes the work and provides ideas for future research.
@@ II-A Running Example
We start by making a running example so that we can describe the Flexoffer concept in the following subsections. Here, we will consider the case of a heat pump operating in a room. The heat pump is a Daikin Altherma [15], with maximum operational power $P_{max}=4.6$ kW and coefficient of performance (COP) $3.65$. This model supports the SG-Ready interface [16, 17], a standardized smart grid communication protocol that enables seamless integration with modern energy management systems. The room has a rectangular section with sides of $4$ and $5$ meters respectively, and it is $3$ meters high. The heat transfer coefficient is $c_{ht}=6\frac{W}{m^{2}K}$ [18], and heat is dispersed through one of the $3$m x $4$m side walls. The room temperature is $295$K ($22^{\circ}$), while the outside temperature is $275$K ($2^{\circ}$). We assume that room temperature must stay between $293$K ($20^{\circ}$) and $297$K ($24^{\circ}$), respectively, because of user settings. For reference, the amount of energy (heat) needed to maintain the temperature to $295$K for one hour can be obtained by
@@ II-B FlexOffer Life Cycle
The Flexoffers model describes flexibility in a unified format for different device types [11, 19], by providing an approximated representation of flexibility. In exchange for the small amount of accuracy lost with respect to exact models, modeling flexibility with FOs allows for much better scalability over long time horizons and high numbers of devices. FOs provide constraints for energy consumption: optimizing an FO means solving an optimization problem on energy consumption (e.g., cost minimization) within the constraints given by the FO. The solution of this optimization problem is a set of values $(e_{1},\ldots,e_{T})$ for energy consumption at each time unit: this is called a schedule. FOs can also be aggregated: this means that, from $M$ FOs, it is possible to generate $N\ll M$ FOs that represent their combined flexibility, with some losses. The opposite process is called disaggregation: from a schedule for an aggregated FOs, it is possible to generate schedules for the original FOs so that their combined energy consumption is the same as the original schedule.
@@
Figure 1: A schematic description of the FOs life cycle (taken from [21]).
@@ II-C Standard FlexOffers
The simplest FO constraints are start time constraints and slice (energy) constraints. We are considering time as discrete and divided into time units, also called time slices: their duration is one hour in the running example. Start time constraints define the earliest and latest time at which the load can start. Energy constraints define, at each time unit, the minimum and maximum amount of energy that can be consumed from the considered energy load: in other words, at each time unit $t$, two numbers $emin_{t}$ and $emax_{t}$ are defined such that $emin_{t}\leq e_{t}\leq emax_{t}$, where $e_{t}$ is the energy consumption during time unit $t$. An FO with just these constraints is called a Standard FlexOffer (SFO). We can see an SFO example in Figure 2(a) for the running example for eight time units. The pink bars indicate the amount of energy that can be consumed at each time unit: here, at each time unit, it varies between 0.303 kWh and 0.478 kWh. At each time, the horizontal bars indicate a possible amount of energy consumption $e_{t}$.
@@ II-F Uncertain FlexOffers
Uncertain FOs (UFOs) [21] are a type of FOs that consider uncertainty from flexibility devices behavior. Two main types of uncertainty are considered: time and amount uncertainty. At each time unit, time uncertainty refers to the probability for the device to be operative at that time, while amount uncertainty refers to the amount of energy that can be consumed by the device at that time. A UFO is created in two steps. First, uncertainty related to the state (in the running example, temperature) is modeled at each time $t$; second, the probability for each energy value at each time to be feasible is calculated, taking into account both types of uncertainty. Some functions $\{f_{1},\ldots,f_{T}\}$ describing those probabilities will be obtained: those functions will define the UFO. UFOs can be visualized by choosing a probability threshold $P_{0}$. At each time $t$, the energy values having probability at least $P_{0}$ of being available for consumption can be described by intervals. Figure 2(d) shows this visualization for the running example: if we choose $P_{0}=1$, at each time unit, the available energy values are described by the pink bars, while if we choose $P_{0}=0.8$, the available energy values are described by the combined pink and blue bars. This is because at the first time unit, the status of the device is known, and the amount of available flexibility can be determined exactly, while for the next time units, there is more uncertainty (represented here by the blue bars) and the amount of flexibility certainly available (pink bars) is lower.
@@ II-G SAREF and SAREF4ENER
Smart Applications REFerence (SAREF) [^2] https://saref.etsi.org/ [12] is an ontology for IoT devices and solutions published by ETSI [^3] https://www.etsi.org/ in a series of Technical Specifications initially released in 2015 and updated in 2017. Even though its initial objective was to build a reference ontology for appliances relevant to energy efficiency, SAREF can also serve as a reference model to enable better integration of data from various vertical domains in the IoT. SAREF has been extended to different domains: specifically, SAREF for Energy Flexibility (SAREF4ENER) [13] is its extension for the energy domain, and aims to provide a standardized way to describe and interoperate energy-related devices, systems, and services. It allows us to describe devices such as solar panels, electric vehicles, and smart meters, as well as systems like microgrids, and energy management/demand response services. By using SAREF4ENER, developers can create applications that easily integrate and communicate with different energy devices, systems, and services, regardless of the manufacturer or protocol used, which enables more efficient and sustainable energy management.
@@
The World Wide Web Consortium (W3C) has developed a standard for enabling Internet of Things (IoT) interoperability, whose name is W3C Web of Things (WoT) [22]. In this context, Thing Descriptions (TDs) [23] have been developed: they describe the metadata and interfaces of the considered devices that participate in the WoT. Things are organized in TD Directories (WoT TDDs), which make it possible to register, manage, and search TDs. An important feature for IoT consumers is to send queries to the WoT directory: this is achieved by having a common ontology to annotate TDs; however, before dCO, no ontologies were integrating WoT into smart buildings. dCO [14] has been created with the purpose of achieving this integration: in this paper, we will focus on the aspect of dCO that strictly concerns energy flexibility. A description of the modules can be found in Appendix -A). The contribution of our work is the energy flexibility module, which is not described in [14]: in this paper, we explain how it extends SAREF4ENER.
@@
This section shows how SAREF4ENER has been extended to describe FOs. We show how we designed the dCO flexibility module and how we encode attributes and constraints. Details on FO messages can be found in Appendix -B.
@@ III-B Energy Flexibility Ontological Design
  • dco:hasEnergyConstraintList: (max 1 s4ener:EnergyMin, max 1 s4ener:EnergyMax). In Figure 2(a), they are 0.303 and 0.478 respectively);
@@
  • dco:totalEnergyConstraint: (max 1 s4ener:EnergyMin, max 1 s4ener:EnergyMax). In Figure 2(b), they are 2.592 and 3.381 respectively);
  • dco:dependencyEnergyConstraintList: Contains the dependency constraints for energy in a Dependency FlexOffer;
  • dco:uncertainFunctions: (min 1 Function): Contains the parameters which define the uncertain FlexOffer;
  • dco:uncertainThreshold: a real number that represents the probability threshold. In the running example in Figure 2(d), it can be 1 (pink) or 0.8 (blue).
@@
  • dco:scheduleID: the id of a schedule;
  • dco:updateID: the id of the schedule update;
@@ IV Evaluation
The domOS project demonstrated dCO feasibility in real-world case studies in Denmark, France, and Switzerland: all functionalities, e.g., querying building descriptions, are operative, and interoperability between smart buildings and services is achieved. The energy services using dCO cut down energy consumption and smoothened load consumption curves for 124 individual customers and 50 days. The cumulative amount of flexibility (load shedding) was several mWh.
Regarding flexibility representation, SAREF4ENER is able at most to accurately represent flexibility for devices that only need this type of constraint. This is the case with wet devices, such as dishwashers and washing machines, as shown in [11]. These devices have a load with a fixed power/amount profile, which is only shifted in time. However, for other devices, SFOs/TECFOs are outclassed by DFOs in terms of accuracy in flexibility representation, as we can see in previous literature. The ability to retain flexibility for a model is calculated by the economic metric [19], defined as the amount of profit that can be obtained by optimizing flexibility with the objective of cost reduction. Specifically, for the charge and discharge process of batteries, DFOs retain $61$% of the flexibility, against the $38$% from TECFOs and $10$% from SFOs [19]. UFOs can perform even better, retaining up to $66.4$% of flexibility, which goes up to $86.8$% of flexibility when imbalance penalties for violating constraints are high (high penalty case) [21]. Regarding EVs, flexibility modeling via DFOs is more accurate than SFOs and TECFOs, as slice constraints and TECs are only a subset of DFO constraints, and therefore the amount of flexibility retained by DFOs is higher: yet, in [21], we see that UFOs can retain $92$% of flexibility, against the $77.3$% retained by DFOs. For the high penalty case, UFOs retain $87.5$% of flexibility, against the $61.8$% from DFOs, which still outperform SFOs and TECFOs. Finally, regarding heat pumps, the comparison between DFOs and SFOs has been done in [24], showing that DFOs can retain more than three times the amount of flexibility from SFOs: in [15] DFOs for heat pumps have been further improved, showing that they can retain up to $98.9$% of flexibility according to the economic metric. The ability for dCO to model the significant amounts of flexibility for these devices ensures that developers can create applications that allow for an accurate description of flexibility for batteries, EVs and heat pumps. Table I summarizes this, showing which use cases are supported by the dCO flexibility module and SAREF4ENER, respectively.
@@
In order to validate dCO, we also used the OntOlogy Pitfall Scanner! (OOPS!) tool [^4] https://oops.linkeddata.es/ to evaluate it. OOPS! [25] is a tool, well-known in the Semantic Web community, designed with the purpose of identifying pitfalls and design errors in ontologies. There are three levels of importance in pitfalls identified by OOPS!. Critical pitfalls need to be resolved, as they may affect the ontology consistency, reasoning, and applicability. Important pitfalls would improve the ontology’s robustness if fixed, but they do not undermine the ontology’s function. Minor pitfalls would improve the ontology’s readability if fixed, but do not pose problems for the ontology. OOPS! identified quite a few minor and important pitfalls, but only 11 critical ones. However, the critical pitfalls were all of the type Defining multiple domains or ranges in properties. In our case, it was not only valid, but also necessary for properties to have multiple domains and ranges to capture the complexity of our flexibility scenario. A simple example of this is the dco:hasEnergyConstraintList property, which can have up to one s4ener:EnergyMin attribute and up to one s4ener:EnergyMax attribute. For this reason, we intentionally designed our ontology with this feature, and therefore none of the critical pitfalls identified by the OOPS! tool undermines the function of the ontology, rather, they are exactly as they should be. Thus, the ontology meets the checked requirements.